OpenStudy (anonymous):

what is a number that can be divided by 3 and have a reminder of 1 , the same number and divided 4 reminder of 1 , same number divided by 5 and reminder of 1 , same number and divided by 6 and reminder of 1 , same number and divided by 7 and reminder of 1?

OpenStudy (anonymous):

The easiest way to find a number is 3*4*5*6*7 + 1 = 2520 + 1 = 2521

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Ryaan

OpenStudy (oaktree):

You should use the Chinese Remainder Theorem.

OpenStudy (anonymous):

i already know what the answer is